Cacodylic acid is a common weak acid used to buffer biochemical reactions and has a pKa of 6.2. If you were to make a 1 L buffer of 100 mM cacodylic acid at pH 6.7, what volume of 5 M NaOH is needed to make this? Show your calculations
Answer #1

let we have V L of NaOH is used
then, number of mole = (molarity)*(volume in L)
number of mole of NaOH = 5V
then, number of mole of cacodylic acid = 0.1 mole
so,
5V mole of NaOH react with 5V mole of acid to form 5V mole of salt
number of mole of acid remain = 0.1-5V

as volume for salt and acid remainig is same
molarity ratio of salt and acid remaing = number of mole of ratio of salt and acid
use,
pH = pKa + log([salt]/[acid])
6.7 = 6.2 + log(5V/(0.1-5V))
log(5V/(0.1-5V)) = 0.5
5V/(0.1-5V) = 10^0.5
5V/(0.1-5V) = 3.16
5V = 3.16*(0.1-5V)
5V = 0.316 - 15.8V
20.8V = 0.316
V = 0.0152 L
= 15.2 mL

Answer : 15.2 mL